# Article Type: Pre-Print

### v1

@@ Introduction
  • Assess impact values: We assign numerical values to each threat based on its impact level on safety, financial stability, operational continuity, and compliance with privacy and legislation. (Table 1)
    Table 1: Mapping of impact levels to numerical values
    Impact Level Value
    None 0
    Low 1
    Medium 10
    High 100
  • Calculate impact score: Equation 1 is used to compute the overall impact score of each threat. This equation combines the weighted values of safety, financial, operational, and privacy and legislation impacts. To reflect the critical importance of safety in the context of automotive cybersecurity, we set $w_{s}=10$, placing the maximum weight on safety.
    Impact $\small\text{Impact Score}=\frac{w_{s}i_{s}+w_{f}i_{f}+w_{o}i_{o}+w_{pl}i_{pl}}{w_{s}+w_{f}+w_{o}+w_{pl}}$ (1)
    where $w_{s}=10$ and $w_{f}=w_{o}=w_{pl}=1$.
  • Derive impact rating: Based on the computed impact score of each threat, we derive the corresponding ratings using the thresholds defined in Table 2. This rating helps prioritise threats based on their potential severity and guides subsequent risk mitigation in our analysis.
    Table 2: Mapping of computed impact scores to ratings
    Impact Score Rating
@@ 0.00 None
The final impact evaluation is tabulated in Figure 4.
Figure 4: Impact assessment based on HEAVENS 2.0
Step 5: Model attacks
@@
Figure 5: Attack tree for IVI system
Step 6: Evaluate attack feasibility
Next, we use CVSS version 3.1 to analyse and compute the attack feasibility ratings. CVSS provides a standardised method for assessing the severity of vulnerabilities based on their impact and exploitability factors.
The CVSS 3.1 base score is computed using Equation 2.
Base Score=Round up(0.6×Impact+0.4×Exploitability1.5Impact $\tiny\text{Base Score}=\text{Round up}\left(\frac{{0.6\times\text{Impact}+0.4\times\text{Exploitability}-1.5}}{{\text{Impact Factor}+0.8}}\right)$ (2)
where Impact, Exploitability, and Impact Factor are derived from the following components:
  • Impact Metrics (C, I, A):
    • Confidentiality (C): The potential impact on confidentiality.
    • Integrity (I): The potential impact on integrity.
    • Availability (A): The potential impact on availability.
  • Exploitability Metrics (AV, AC, PR, UI):
@@
Additionally, the CVSS temporal score takes into account factors such as exploit code maturity, remediation level, and report confidence. It is calculated using Equation 3.
Temporal Score=Round up(Base Score×Exploit Code Maturity×Remediation Level×Report $\tiny\begin{split}\text{Temporal Score}=\text{Round up}(\text{Base Score}\times\text{Exploit Code Maturity}\times\\ \text{Remediation Level}\times\text{Report Confidence})\end{split}$ (3)
Figure 6: Attack feasibility ratings based on CVSS version 3.1
Figure 6 lists the attack feasibility ratings computed using the CVSS version 3.1 calculator (https://nvd.nist.gov/vuln-metrics/cvss/v3-calculator). These ratings are based on the information available at the time of assessment during the early design stage of the IVI system.
Figure 7: Risk ratings evaluated using a 5x5 risk matrix
Step 7: Evaluate risks
In this step, we evaluate the cybersecurity risks of the IVI system using a 5x5 risk matrix. The corresponding risk ratings for identified threats are tabulated in Figure 7.
These risk ratings serve as a guideline for the Original Equipment Manufacturer (OEM) of the vehicle or the manufacturer of the IVI system. They can use these ratings to prioritise the implementation of mitigation controls during the design phase of the system life cycle. This approach ensures that adequate attention and resources are allocated to address the most critical cybersecurity concerns identified through this comprehensive analysis.
Step 8: Review, update, and iterate
@@
Updating the Security Model After a Vulnerability Disclosure
In the earlier section, we discussed the TARA framework as part of our approach to security modelling for automotive systems. Generally, this is performed during the early stages of system life cycles. However, evolving cybersecurity threats necessitate the incorporation of new knowledge gained from ongoing security monitoring and threat intelligence throughout the system’s life cycle.
ACTISM can be seamlessly integrated with the TARA framework to create a threat-informed dynamic security modelling methodology.
We now apply this methodology to a real-world example using the Tesla Electric Vehicle (EV) jail-break vulnerability disclosure [13].
@@
The updated risk rating based on these revised metrics is shown in Figure 8. In summary, the corresponding risk changed from Medium to High.
This ability to perform timely updates to the risk assessment is crucial for maintaining robust cybersecurity in automotive systems. By continuously adapting to new threats and changes in attack feasibility and impact, proactive measures can be implemented to safeguard automotive system safety and functionality. This dynamic approach not only enhances the resilience of automotive systems against evolving cybersecurity threats but also reinforces the importance of informed decision-making in the face of emerging risks. Ultimately, it empowers stakeholders to take appropriate actions that protect both the automotive system and its users.
Figure 8: Updated risk rating following vulnerability disclosure
Discussion
Unlike early-stage security modelling, where OEMs and IVI system manufacturers implement comprehensive measures, addressing cyber-physical vulnerabilities later in the life cycle often involves temporary fixes or workarounds. These may not fully mitigate the risks, so vehicle owners can use risk ratings to make informed decisions. By adopting additional precautions, they can enhance the cybersecurity resilience of their vehicles throughout their life cycle.
This underscores the need for a continuous, iterative, and threat-informed security modelling process. Ongoing threat intelligence gathering, including real-time monitoring and tracking vulnerability disclosures, is crucial. A comprehensive repository of vulnerability attributes, stored in a standardised format such as JSON, enables automation. Tools can parse this data, extract relevant information, and enrich the security model via workflows like TARA.
CVSS scores and risk ratings could then be dynamically recalculated based on updated threat context. Security models can be formalised using Domain-Specific Languages (DSLs) such as the Meta Attack Language (MAL), allowing clearer descriptions of threats and vulnerabilities. Furthermore, adversary tactics from the MITRE ATT&CK knowledge base could be incorporated to enhance these models.
Machine learning approaches could also help prioritise vulnerabilities by analysing their impact, using historical data and threat intelligence to improve automotive cybersecurity.
Conclusion
Automotive systems are increasingly facing cybersecurity threats, with attackers targeting vulnerabilities that pose risks to data, functionality, and safety. Ensuring the security of automotive systems is critical to maintaining their integrity and reliability.
@@ Conclusion
We introduced ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling), a novel approach designed to enhance the resilience of automotive systems by dynamically updating their cybersecurity posture. ACTISM responds to evolving threats and attacker tactics, ensuring security models remain relevant throughout a system’s lifecycle.
Unlike traditional models, ACTISM continuously integrates new data from threat intelligence and security monitoring, keeping protection up to date. It also captures the multi-agent nature of cyber-physical attacks, making it applicable to automotive domains. This integration strengthens the accuracy and completeness of security models, reflecting the complexity of real-world cyber-physical attacks.
Cybersecurity incidents in automotive systems differ from traditional IT breaches due to their potential physical consequences, such as vehicle misnavigation or compromised safety. ACTISM’s consequence-driven approach ensures that cyber-physical risks are considered, directly addressing the operational and safety integrity of automotive systems.
@@
To demonstrate ACTISM’s practical application, we developed a comprehensive security model for an IVI system. This involved identifying key assets, conducting threat modelling, and performing risk assessments, resulting in a robust security framework. A real-world case study further highlighted ACTISM’s ability to adapt to emerging threats and its potential for broader use in other automotive technologies.
In conclusion, ACTISM offers a proactive, adaptable approach to automotive cybersecurity, addressing current challenges while preparing for the evolving landscape of automotive technology.

### v2

@@ 4.4 Step 4: Evaluate impact
  • Assess impact level: By definition in the HEAVENS framework, the impact level of a specific threat is an estimate of the expected loss for various stakeholders when the threat is realised. To estimate the impact, we use four parameters directly related to the security objectives defined in HEAVENS: safety (s), financial (f), operational (o), and privacy (p).
    • Safety Impact: Refers to the safety of vehicle occupants, road users, and infrastructure.
    • Financial Impact: Encompasses all direct and indirect financial damages incurred by stakeholders.
    • Operational Impact: Pertains to operational damages that have minimal or no safety or financial consequences. For example, the loss of secondary functionalities such as entertainment systems.
    • Privacy Impact: Deals with damages resulting from privacy violations.
    Each of the four parameters is assigned one of four impact levels: None, Low, Medium, or High. For the purposes of this experience report, we used the impact level definitions from [12]. Impact levels are typically assigned by automotive domain experts who have detailed knowledge of the automotive system, its vulnerabilities, and corresponding impacts on security objectives. In our running example, we assign impact levels based on our assessment of potential safety, operational, financial, and privacy consequences. This assessment is based on our understanding from the literature, and may differ from evaluations of experts in the field.
    Next, numerical values are assigned to the impact levels. These values are chosen in a logarithmic scale (0, 1, 10, 100) to effectively capture the exponential increase in the impact levels. This helps in clearly differentiating between the levels of impact related to the security objectives, and ensures that high-impact threats are given appropriate priority.
    Table 10 [8] categorises the impact levels and parameter values according to the security objectives.
  • Calculate impact score: After estimating the individual impact levels of the security objectives and their corresponding parameter values, the overall impact score is computed using the following equation:
    $\small I=w_{s}i_{s}+w_{o}i_{o}+w_{f}i_{f}+w_{p}i_{p}$ (1)
    Normalisation is necessary to ensure that the overall impact score is on a consistent scale, allowing for fair and meaningful comparisons across different systems and contexts. This uniformity is crucial when comparing the cybersecurity posture of various systems, such as between automotive systems.
    The overall normalised impact score is computed using the following equation:
    $I_{\text{norm}}=\frac{w_{s}i_{s}+w_{o}i_{o}+w_{f}i_{f}+w_{p}i_{p}}{100*(w_{s}+w_{o}+w_{f}+w_{p})}$ (2)
    Weight values $w_{s}$, $w_{o}$, $w_{f}$, and $w_{p}$ are selected based on the specific needs of the use case being evaluated. In the case of an automotive system, the safety impact is the most severe consequence for stakeholders, such as vehicle occupants, road users, and infrastructure. Comparatively, financial, operational, and privacy impacts may have lower consequences. As such, the weights are selected as: $w_{s}=10$ and $w_{o}=w_{f}=w_{p}=1$. Equations 1 and 2 are simplified as follows:
    $I=10i_{s}+i_{o}+i_{f}+i_{p}$ (3)
    $\small I_{\text{norm}}=\frac{10i_{s}+i_{o}+i_{f}+i_{p}}{1300}$ (4)
  • Assign impact rating: Each threat is assigned an impact rating based on the impact score $I_{\text{norm}}$ of each assessed threat, matched against thresholds tabulated in Table 2, as referenced from [12].
    Table 2: Impact rating table
    Impact Score Impact Rating
@@ 0.00 None
For illustration, we show how the impact rating of Threat ID 9146 from Table 9 [8] is derived using this methodology. In this specific threat, an adversary may reverse engineer the head unit firmware to find sensitive information, resulting in information disclosure from the IVI System’s local maps data store. Using the definitions of the impact assessment table (Table 10 [8]), we evaluate $i_{s}=0$, $i_{o}=0$, and $i_{f}=0$, as there are no injuries or discernible operational and financial effects resulting from the disclosure of information. However, disclosure of personally identifiable information or passwords from the data store can result in privacy violations leading to abuses such as impersonation of the victim to perform actions with stolen identities. As such, the impact level for privacy is assessed as Medium and $i_{p}=10$. Using Equation 2, $I_{\text{norm}}\ =0.0077$, which corresponds to a Low impact rating.
Table 3: Impact rating example
Threat ID $i_{s}$ $i_{o}$ $i_{f}$ $i_{p}$ Impact Score Impact Rating
9146 0 0 0 10 0.0077 Low
4.5 Step 5: Model attacks
@@
Figure 3: Attack tree for In-Vehicle Infotainment (IVI) system
4.6 Step 6: Evaluate attack feasibility
Next, we use CVSS version 3.1 to analyse and compute attack feasibility. CVSS provides a standardised method for assessing the severity of vulnerabilities based on their impact and exploitability factors, as described below.
  • Impact Metrics:
    • Confidentiality (C): Impact on the confidentiality of data, such as driver information, vehicle location, and system diagnostics.
    • Integrity (I): Impact on the integrity of automotive systems, including the accuracy and reliability of data and control signals within the vehicle.
    • Availability (A): Impact on the availability of automotive functions and services, such as the ability to start the vehicle or use navigation systems.
  • Exploitability Metrics:
@@ 4.6 Step 6: Evaluate attack feasibility
In this experience report, instead of using the CVSS base score, which is common in the literature, we use the CVSS temporal score to incorporate and emphasise the temporal aspects of the vulnerability, such as exploit code maturity, remediation level, and report confidence. These are shown at Equations 5 and 6.
CVSS Base Score=Round up(0.6×Impact+0.4×Exploitability1.5Impact $\text{CVSS Base Score}=\text{Round up}\left(\frac{{0.6\times\text{Impact}+0.4\times\text{Exploitability}-1.5}}{{\text{Impact Factor}+0.8}}\right)$ (5)
CVSS Temporal Score=Round up(CVSS Base Score×Exploit Code Maturity×Remediation Level×Report $\begin{split}\text{CVSS Temporal Score}=\text{Round up}(\text{CVSS Base Score}\times\text{Exploit Code Maturity}\times\\ \text{Remediation Level}\times\text{Report Confidence})\end{split}$ (6)
Next, the CVSS temporal scores for the identified threats are computed using the CVSS version 3.1 calculator (https://nvd.nist.gov/vuln-metrics/cvss/v3-calculator), then mapped to qualitative severity ratings using Table 4 as referenced from the CVSS 3.1 Specification Document. The severity ratings, which we refer to as attack feasibility in this experience report, are tabulated in Table 5.
Table 4: Qualitative severity rating scale
CVSS Score Severity Rating (Attack Feasibility)
0.0 - 3.9 Low
4.0 - 6.9 Medium
7.0 - 8.9 High
9.0 - 10.0 Critical
Table 5: Attack feasibility based on CVSS version 3.1
Threat ID AV AC PR UI S C I A E RL RC CVSS Score Attack Feasibility
9158 A L N N C N L H P U U 7.1 High
9132 A H N N C N H H U U U 6.7 Medium
9089 P L N R U H H H U U U 5.6 Medium
9144, 9060 A L L R U H H H U U U 6.2 Medium
9146 P H N R U H L N U U U 3.9 Low
4.7 Step 7: Evaluate risks
In this step, we evaluate the cybersecurity risks of the IVI system using a 5x5 risk matrix illustrated in Figure 4. In the context of automotive impact assessment, where the stakes can be high and risks multifaceted, a 5x5 risk matrix provides the depth and granularity needed to accurately assess and manage risks, compared to a simpler 3x3 risk matrix. The risk ratings for the identified threats are tabulated in Table 6.
Figure 4: 5x5 risk matrix
Table 6: Risk ratings based on attack feasibility and impact ratings
Threat ID Description Attack Method Source Attack Feasibility Impact Rating Risk Rating
9132 Cause a denial-of-service of the central gateway, affecting vehicle functions Inject or flood invalid data to the central gateway causing it to crash or stop. IVI System High Critical Extreme
9159 Cause gateway to misbehave, crash or stop, affecting vehicle functions Exploit vulnerabilities, through outdated software or configuration weaknesses. IVI System Medium High High
9089 Elevate privileges on the IVI System (via USB) Exploit vulnerabilities, through outdated software or configuration weaknesses, via USB removable media. USB Medium Low Medium
9144, 9060 Elevate privileges on the IVI System (via Bluetooth or Wi-Fi) Exploit vulnerabilities, through outdated software or configuration weaknesses, via Bluetooth or Wi-Fi. Bluetooth or Wi-Fi Medium Low Medium
9146 Access sensitive data e.g. passwords from IVI System Embedded Data Store Reverse engineer the head unit firmware to find keys to access the IVI System Embedded Data Store. Physical Access Low Low Medium
These risk ratings serve as a guideline for the Original Equipment Manufacturer (OEM) of the vehicle or the manufacturer of the IVI system. They can use these ratings to prioritise the implementation of mitigation controls throughout the system life cycle. This approach ensures that adequate attention and resources are allocated to address the most critical cybersecurity concerns identified through this comprehensive analysis.
4.8 Step 8: Review, update, and iterate
@@
5 Updating the Security Model After a Vulnerability Disclosure
ACTISM enhances security modelling by introducing a dynamic and iterative methodology, resulting in a threat-informed, evolving security model. This approach addresses evolving cybersecurity threats that necessitate the incorporation of new knowledge gained from security monitoring and threat intelligence throughout the system’s life cycle.
We now apply this methodology to a real-world example using the Tesla Electric Vehicle (EV) jail-break vulnerability disclosure [14]. Although this vulnerability is specific to the Tesla EV platform, ACTISM is designed to be applicable to all automotive systems (including non-EVs) and their relevant vulnerabilities.
@@ 5 Updating the Security Model After a Vulnerability Disclosure
With these updated parameters, the attack feasibility increased from 2 (Low) to 3 (Medium), as shown in Table 7.
Table 7: Updated attack feasibility following vulnerability disclosure
Threat ID AV AC PR UI S C I A E RL RC CVSS Score Attack Feasibility
9158 A L N N C N L H P U U 7.1 High
9132 A H N N C N H H U U U 6.7 Medium
9089 P L N R U H H H U U U 5.6 Medium
9144, 9060 A L L R U H H H U U U 6.2 Medium
9146 P H N R C H H H P T C 6.4 Medium
Besides attack feasibility, the impact assessment of Threat ID 9146 in the security model also needs updating. Using the definitions from the impact assessment table (Table 10 [8]), we estimate the impact on safety to increase from None to Low ($i_{s}=10$), as there is now a possibility of light and moderate injuries, should the attacker access and manipulate certain vehicle functions via the vulnerability. We also expect the operational impact to increase from None to Low ($i_{o}=10$), as there may be a degradation of primary vehicle functions. The financial impact is also expected to increase from None to Low ($i_{f}=10$), as the financial damage from injury claims may be substantial but not threatening the existence of the manufacturer. Lastly, the privacy impact is expected to increase from Low to High ($i_{p}=100$), as the privacy violation of multiple vehicle owners and drivers may lead to extensive media coverage and severe loss of market share. Using Equation 2 to compute the updated normalised impact score, $I_{\text{norm}}\ =0.1692$, which corresponds to a High impact rating (Table 8).
Table 8: Updated impact rating following vulnerability disclosure
Threat ID $i_{s}$ $i_{o}$ $i_{f}$ $i_{p}$ Updated Impact Score Updated Impact Rating
9146 10 10 10 100 0.1692 High
With the updated attack feasibility (Medium) and impact rating (High), the overall risk rating for Threat ID 9146 increased from Medium to High, in accordance with the 5x5 risk matrix (Figure 4). This ability to perform timely updates to the risk assessment is crucial for maintaining robust cybersecurity in automotive systems. Ultimately, it empowers stakeholders to adopt additional precautions and improve the cybersecurity resilience of their vehicles throughout their life cycle.
6 Future directions
While ACTISM provides an approach to dynamically updating security models with both cyber and physical threats, it is not without its limitations. One of the primary challenges is how the accuracy of the model can be validated given the constantly evolving nature of threats and the potential for incomplete or inaccurate threat intelligence data. Moreover, while ACTISM’s adaptability to new threats is a strength, it can also pose a challenge in maintaining the models’ integrity and consistency over time. Despite these limitations, the potential future directions outlined below can address these challenges and enhance the robustness of ACTISM.
Ongoing threat intelligence gathering, including real-time monitoring and tracking of vulnerability disclosures, is crucial for keeping the security model current. Establishing a comprehensive repository of vulnerability attributes, stored in a standardised format such as XML or JSON, can enable automation and streamline the updating process. Additionally, formalising security models using Domain-Specific Languages (DSLs) like Automation Markup Language (AutomationML) and Meta Attack Language (MAL) can provide clearer descriptions of threats and vulnerabilities.
Incorporating adversary tactics from the MITRE ATT&CK knowledge base can further enhance these models by offering insights into real-world attack strategies. Moreover, leveraging machine learning approaches to prioritise vulnerabilities based on their impact, utilising historical data and threat intelligence, can significantly improve automotive cybersecurity.
These potential future directions can ensure that automotive cybersecurity remains resilient and adaptive to emerging cybersecurity challenges.
7 Conclusion
The cybersecurity landscape for automotive systems is rapidly evolving, with attackers increasingly targeting vulnerabilities that threaten data integrity, functionality, and safety. Traditional static security assessment methodologies, such as STRIDE, often fall short in addressing the dynamic and multi-faceted nature of these threats. There is a clear knowledge gap in the ability to model and respond to security threats in a dynamic, iterative manner tailored specifically for automotive systems.
@@ 7 Conclusion
In response, we introduced ACTISM (Automotive Consequence-Driven and Threat-Informed Security Modelling), an integrated approach designed to enhance the resilience of automotive systems by dynamically updating their cybersecurity posture. ACTISM effectively responds to evolving cyber-physical threats and attacker tactics, ensuring that security models remain robust and relevant throughout a system’s lifecycle.
@@
To illustrate the practical application of ACTISM, we developed a comprehensive security model for an IVI system. This process included identifying key assets, conducting thorough threat modelling, and performing risk assessments, ultimately resulting in a robust security framework. A real-world case study involving Tesla’s IVI system further demonstrated ACTISM’s ability to adapt to emerging threats and underscored its potential for broader applicability across various automotive technologies.
In conclusion, ACTISM addresses the critical knowledge gap in automotive cybersecurity by providing a proactive, adaptable approach to security modelling. This approach not only tackles current challenges but also prepares for the future landscape of automotive technology.